If A SKI RACER TRAVELS 2.03 MILES IN 2 MINUTES AND 57 SECONDS. WHAT IS HIS AVERAGE SPEED?
Found 4 solutions by checkley71, misscrt, stanbon, guadalupeh_14:
Answer by checkley71(8403)

2.03/(2 57/60)=2.03/[(120+57)/60]=2.03/[(177)/60]=2.03/2.95=.6881355 MPS. ANSWER.

Answer by misscrt(37)

If A SKI RACER TRAVELS 2.03 MILES IN 2 MINUTES AND 57 SECONDS. WHAT IS HIS AVERAGE SPEED?
Total time traveled is equal to the sum of 2 minutes and 57 seconds.
Convert 2 minutes to hours and 57 seconds to hours. Then add the two times.
Then, Velocity = Distance/Time
[Note: The distance given is 2.03 miles.]
T1 = 2 min/(60 min) * 1 hour = .033 hr
T2 = 57 sec/(60 sec)*(1 min)/(60 min)* 1 hour = .016 hr
Total time, T, = T1 + T2 = .033 + .016 = .049
Velocity = 2.03 miles/(.049 hr)
= 41.43 miles per hour
Hope that helps.
Regards,
misscrt

Answer by stanbon(48568)

If A SKI RACER TRAVELS 2.03 MILES IN 2 MINUTES AND 57 SECONDS. WHAT IS HIS AVERAGE SPEED?
------------------
rate = distance/time = 2.03 miles / 2 (57/60) minutes
= 2.03 miles/2.95 minutes
= 0.6881 miles per minute
========================
Cheers,
Stan H.

Answer by guadalupeh_14(1)